in a triangle ABC the median to side BC is of the length 1/root11-6root3 and it divides the angle into angles of 30degree and 45degree find the length of side BC

Grade:11

2 Answers

Vijay Mukati
8 years ago
Dear student,

In triangle ABC, if AD is median and a is length of side BC, then
Using sin Rule:
SinB = AD*sin (30)/(a/2)
sinC = AD*sin(45)/(a/2)
Now add the two above equation and apply the formula for SinC + sinB and convert it into A by using (A + B + C = pi).
And simplify it.

Thanks.
siva prasad reddy
8 years ago
In triangle ABC, if AD is median and a is length of side BC, thenUsing sin Rule:SinB = AD*sin (30)/(a/2)sinC = AD*sin(45)/(a/2)Now add the two above equation and apply the formula for SinC + sinB and convert it into A by using (A + B + C = pi).And simplify it.